OpenStudy (falling_in_katt):

Which statement is true? A. (p2)5= p3 B. (p2)5= p7 C. (p2)5= p10 D. (p2)5= p25

OpenStudy (anonymous):

\[\left( x^a \right)^b=x ^{a*b}\] \[\left( p^2 \right)^5=?\]

OpenStudy (falling_in_katt):

C